# Schwinger dark matter production

### v1

@@ III Cosmic evolution and dark
Above $\overline{T}_{\chi}$ the energy density of $\chi$ evolves according to Eq. (41). When $\chi$ becomes non-relativistic the Schwinger production is no longer effective. We must also ensure that $\rho_{\chi}(\overline{T}_{\chi})\ll\rho_{A}(\overline{T}_{\chi})$ is satisfied all the way down to $\overline{T}_{\chi}$ in order to justify neglecting the back reaction of the Schwinger production on the dark electric field as we have assumed in our analysis. This is always ensured when $g_{D}$ satisfies Eq. (49) so our analysis is self consistent. Below $\overline{T}_{\chi}$ the charged particles redshift like matter,
$\rho_{\chi}(T)=\rho_{\chi}(T_{0})\left(\frac{T}{T_{0}}\right)^{3}\,,$ (52)
analogously to Eq. (39). The energy density today is found by equating Eq. (41) and Eq. (52) at $T=\overline{T}_{\chi}$,
$\rho_{\chi}(T_{0})=$ $\frac{1}{2}\bar{\sigma}_{\rm RH}\rho_{E}^{\rm RH}\left(\frac{T_{0}}{T_{\rm RH}}\right)^{3}$ (53)
×$\times\frac{1}{2-\alpha}\left[4\left(\frac{m_{\chi}}{H_{\rm end}}\right)^{\alpha-1}-(2+\alpha)\left(\frac{m_{\chi}}{H_{\rm end}}\right)\right]\,,$
where we have expanded in powers of the small dark gauge coupling keeping only the leading order term and recalling that $\bar{\sigma}_{\rm RH}\propto g_{D}^{2}$ (see Eq. (26)). This result has several interesting features. If the conductivity remains constant after reheating ($\alpha=0$) the final relic density $\rho_{\chi}$ gets enhanced by $H_{\rm end}/m_{\chi}\gg 1$,
$\rho_{\chi}(T_{0})\simeq\bar{\sigma}_{\rm RH}\rho_{E}^{\rm RH}\left(\frac{T_{0}}{T_{\rm RH}}\right)^{3}\frac{H_{\rm end}}{m_{\chi}}\quad(\alpha=0)\,.$ (54)
Note in this expression $H_{\rm end}/m_{\chi}$ cannot be arbitrarily large because it is bound by Eq. (50). An enhancement of $(H_{\rm end}/m_{\chi})^{1-\alpha}$ persists as long as the conductivity redshifts slowly enough ($10\leq\alpha<$ 1). When $\alpha=1$ we reach a value independent of $m_{\chi}$ to a good approximation,
$\rho_{\chi}(T_{0})\simeq 2\bar{\sigma}_{\rm RH}\rho_{E}^{\rm RH}\left(\frac{T_{0}}{T_{\rm RH}}\right)^{3}\quad(\alpha=1)\,.$ (55)
For values $21<\alpha<$ 2, the final relic abundance is suppressed by $(m_{\chi}/H_{\rm end})^{\alpha-1}$. When the conductivity redshifts like Hubble ($\alpha=2$) in radiation domination,
$\rho_{\chi}(T_{0})=$ $\frac{1}{2}\bar{\sigma}_{\rm RH}\rho_{E}^{\rm RH}\left(\frac{T_{0}}{T_{\rm RH}}\right)^{3}$
×$\times\frac{m_{\chi}}{H_{\rm end}}\left[1-\ln\left(\frac{m_{\chi}}{H_{\rm end}}\right)^{4}\right]\quad(\alpha=2)\,.$ (56)
When the conductivity redshifts faster ($\alpha\gg 2$) this gives,
$\rho_{\chi}(T_{0})=\frac{1}{2}\bar{\sigma}_{\rm RH}\rho_{E}^{\rm RH}\left(\frac{T_{0}}{T_{\rm RH}}\right)^{3}\frac{m_{\chi}}{H_{\rm end}}\quad(\alpha\gg 2)\,.$ (57)
In this last case the contribution to $\rho_{\chi}$ comes entirely from the Schwinger production at the end of inflation with the production after reheating being negligible.
The analysis presented is sensible only if the dark sector does not thermalize. Thermalization with the visible sector is easily avoided by taking the dark sector to be secluded. This implies that a possible kinetic mixing between visible and dark $U(1)$ has to be small. We also need to ensure that the dark sector does not thermalize within itself either. This requires that the rate of interactions between the dark charged particles and the dark photon is always slower than Hubble which we can estimate as follows. As long as $\chi$ is relativistic its number density is given by $n_{\chi}(T)\simeq\rho_{\chi}(T)/q_{\chi}(T)$ and the two-to-two cross section, mediated by the exchange of a dark photon, can be estimated as $\sigma_{2\to 2}\sim g_{D}^{4}/q_{\chi}(T)^{2}$. Here, $\rho_{\chi}(T)$ and $q_{\chi}(T)$ are given by Eq. (41) and Eq. (47) respectively. The scattering rate is $\Gamma=n_{\chi}(T)\sigma_{2\to 2}$ and is found to have two terms at leading order in $g_{D}$, one which scales like $T$ and the other like $T^{\alpha-1}$. So $\Gamma$ decreases more slowly than $H\propto T^{2}$ when $\chi$ is relativistic. The trend changes when $\chi$ becomes non-relativistic at which point $\Gamma$ scales like $T^{3}$. Hence, we must require that $g_{D}$ is small enough that $\Gamma<H$ is satisfied all the way down to the temperature $\overline{T}_{\chi}$. This amounts to the condition,
$g_{D}<$ $10\times 2^{1/6}\frac{\xi^{1/2}}{e^{\pi\xi/3}}\epsilon_{R}^{1/3}(2-\alpha)^{1/6}\sqrt{\frac{m_{\chi}}{H_{\rm end}}}$
×$\times\left[4\left(\frac{m_{\chi}}{H_{\rm end}}\right)^{\alpha}-(2+\alpha)\left(\frac{m_{\chi}}{H_{\rm end}}\right)^{2}\right]^{-1/6}\,.$ (58)
As can be seen, this is a weaker bound for any $\alpha$ compared to Eq. (49) which follows from requiring $\chi$ is non-relativistic by matter radiation equality. Thus, we are guaranteed that the dark sector does not thermalize in the mechanism proposed in this study.
@@
With the various energy densities today in hand we can examine the parameter space in which the observed dark matter relic abundance is reproduced. Taking the observed energy density of cold dark matter today Ade et al. (2016) to be $\rho_{\rm CDM}=9.6\times 10^{-48}\ {\rm GeV}^{4}$, the contribution from the dark vector is given by Bastero-Gil et al. (2019, 2022),
$\frac{\Omega_{A}}{\Omega_{\text{CDM}}}$ $=$ $\frac{\rho_{A}(T_{0})}{\rho_{\text{CDM}}}\simeq$ (59)
$0.2$$\times\left(\frac{m_{A}}{\text{GeV}}\right)\left(\frac{H_{\rm end}}{10^{14}\text{GeV}}\right)^{3/2}\frac{e^{2\pi\xi_{end}}}{\epsilon_{R}^{3}\xi_{end}^{3}}\,.$
The contribution from the dark charged particles is,
$\frac{\Omega_{\chi}}{\Omega_{\text{CDM}}}$ $=$ $\frac{\rho_{\chi}(T_{0})}{\rho_{\text{CDM}}}\simeq$
$10^{10}$ $\bar{\sigma}_{\rm RH}\left(\frac{H_{\rm end}}{10^{14}\text{GeV}}\right)^{5/2}\frac{e^{2\pi\xi_{end}}}{\epsilon_{R}^{3}\xi_{end}^{3}}$
$\times$ $\frac{1}{2-\alpha}\left[4\left(\frac{m_{\chi}}{H_{\rm end}}\right)^{\alpha-1}-(2+\alpha)\left(\frac{m_{\chi}}{H_{\rm end}}\right)\right]\,.$
This is proportional to the value of the conductivity at reheating which is the same as at the end of inflation and proportional to $g_{D}^{2}$ (see Eq. (26)). Clearly, if we switch off the gauge interactions in the dark sector we turn off the dark Schwinger production and we have no relics in the form of dark charged particles. To arrive at Eq. (III) we assumed that the conductivity redshifts as $a^{-\alpha}$ during the radiation dominated era (see Eq. (11)) giving,
$\bar{\sigma}(T)=\bar{\sigma}_{\rm RH}\left(\frac{T}{T_{\rm RH}}\right)^{\alpha}\simeq 10^{-2}g_{D}^{2}\left(\frac{T}{T_{\rm RH}}\right)^{\alpha}\,.$ (61)
In the last equality we used Eq. (26) as an estimate for the relation between the conductivity and the gauge coupling. This holds approximately for both dark charged fermions and scalars Kobayashi and Afshordi (2014); Hayashinaka et al. (2016); Banyeres et al. (2018) and will be useful to explore the relic density parameter space as a function of $g_{D}$. In principle, we can have a mixture of dark vector and dark scalars/fermions as dark matter. The case in which $\Omega_{A}$ dominates reduces to what was studied in Bastero-Gil et al. (2019, 2022). Here we focus on the case in which $\Omega_{\chi}$ dominates. In particular, as discussed previously, the mechanism proposed here is also viable when the dark photon is massless in which case $\Omega_{A}$ is negligible and the dark matter is composed entirely of Schwinger produced dark charged particles.
III.1 Dark matter parameter space
Having followed the cosmic evolution of the dark sector energy density and found an estimate for the relic abundance today, we can go on to estimate the regions of parameter space for a viable dark matter sector. To find these regions, we must ensure that not only the observed cold dark matter relic abundance is reproduced, but also that a number of constraints on the model parameters are satisfied for consistency of the mechanism. To be generic, we briefly summarize the constraints both for the dark photon Bastero-Gil et al. (2019, 2022) and for the dark scalar/fermion.
The first constraint is an upper bound on the dark vector mass which results from requiring $m_{A}<H_{\rm end}$ for efficient tachyonic production during inflation. If a particular species dominates the dark matter relic abundance then its mass is bounded from below by the condition that it becomes non-relativistic, thus behaving like cold dark matter, before matter-radiation equality (MRE). There are also robust constraints coming from astrophysical probes which in the case of fermions can lead to a stronger and absolute lower bound of $\mathcal{O}(100\,eV)$ Di Paolo et al. (2018); Alvey et al. (2021). These lower bounds only apply to the components of the dark matter which contribute sizeably to the relic abundance. Thus we can consider much lighter dark sector particles as long as they only make up a tiny fraction of the dark matter by the time of MRE. We must also require the hierarchy in energy densities at reheating in Eq. (34). Imposing that the visible radiation has energy density lower than that at the end of inflation ($\rho_{R}\ll\rho^{\rm{end}}_{I}$) imposes $\epsilon_{R}<1$. To ensure the Universe is radiation dominated until MRE requires that $\rho_{R}>\rho^{\rm{end}}_{E}$ and translates into an upper bound on $H_{\rm end}$,
$H_{\rm end}<\sqrt{3}\times 10^{2}\xi_{\rm end}^{3/2}e^{-\pi\xi_{\rm end}}\epsilon_{R}^{2}M_{\rm Pl}\,,$ (62)
once we fix $\xi_{\rm end}$ and $\epsilon_{R}$. Finally, we must satisfy the bound on the gauge coupling in Eq. (49), along with the condition in Eq. (50), for $\chi$ to become non-relativistic before MRE. There are other constraints on this mechanism which should be satisfied Bastero-Gil et al. (2019, 2022), but they are weaker than the ones discussed here.
We show contours of $\Omega_{\chi}/\Omega_{\rm{CDM}}=1$ given in Eq. (III) as a function of $m_{\chi}$ vs. $H_{\rm end}$ assuming different scalings ($\alpha=0,1,2,\gg 2$) for how the conductivity redshifts after reheating (see Eq. (11)). The lower contours show the case of $\epsilon_{R}=0.75,$ with $\xi_{\rm end}=1$ (blue solid) and $\xi_{\rm end}=9$ (green dashed) while the upper contours are shown for $\epsilon_{R}=10^{-4}$ with $\xi_{\rm end}=1$ (solid red). We estimate the dimensionless conductivity at reheating by $\overline{\sigma}_{\rm{RH}}=g_{D}^{2}10^{-2}$ (see Eq. (26)) where the dark gauge coupling is taken to be the maximum value allowed by the constraint that $\chi$ becomes non-relativistic by the time of MRE and which depends on $\xi_{\rm end},\epsilon_{R},m_{\chi}$ (see Eq. (49)). Note that once we take $g_{D}$ to be its maximum value the relic abundance no longer depends explicitly on $\xi_{\rm end}$. The dark vector, which can even be massless, is taken to be light enough ($m_{A}\lesssim 10^{-20}\,GeV$) such that it contributes negligibly to the dark sector relic abundance. Thus, in this case the dark matter is composed entirely of Schwinger pair produced particles with (feeble) self interactions mediated by the exchange of ultra light dark photons.
[…]
We see this mechanism generically requires the Hubble scale at the end of inflation to be $H_{\rm end}\gtrsim 10^{8}\,GeV$. We can also see that smaller $\epsilon_{R}$, or lower reheat temperatures, pushes us to larger masses as well as larger $H_{\rm end}$ since we require $m_{\chi}<H_{\rm end}$. This is because smaller $\epsilon_{R}$ pushes us to smaller dark gauge couplings (see Eq. (49)) and thus smaller conductivities which requires larger $m_{\chi}$ to compensate for the additional suppression of $\rho_{\chi}$. For $\xi_{\rm end}=1$ and $\epsilon_{R}=10^{-4}$ the contours (red solid) are cutoff at $H_{\rm end}\sim 10^{11}\,GeV$ due to the requirement that $\rho_{R}>\rho^{\rm{end}}_{E}$ as given by Eq. (62) and which generically rules out smaller values of $\epsilon_{R}$. The maximum dark gauge coupling allowed by Eq. (49) takes on values in the range $g_{D}\sim 10^{-15}-10^{-22}$.  The parameter point $\xi_{\rm end}=9$ and $\epsilon_{R}=10^{-4}$ on the other hand is ruled out by this constraint. For $\xi_{\rm end}=1$ and $\epsilon_{R}=0.75$ (blue solid) we can consider $H_{\rm end}$ all the way up to the cosmological limit of $H_{\rm end}\sim 10^{14}\,GeV$ Ade et al. (2021); Tristram et al. (2022). In this case the maximum allowed dark gauge coupling takes on values in the range $g_{D}\sim 0.01-10^{-15}$. For $\xi_{\rm end}=9$ and $\epsilon_{R}=0.75$ (dashed green) the requirement that the gauge coupling be larger than that allowed by the weak gravity conjecture limits us to the narrow range $H_{\rm end}\sim(10^{8}-10^{9})\,GeV$ and the dark gauge coupling is fixed to be $g_{D}\sim 10^{-10}$.
We see the observed dark matter relic abundance can be obtained for dark charged particles as light as $m_{\chi}\sim 10MeV$ for $H_{\rm end}\sim 10^{14}\,GeV$ and $\xi_{\rm end}\sim 1$ if the conductivity remains constant after reheating ($\alpha=0$). In this case Schwinger production takes place almost entirely after inflation. For $\xi_{\rm end}=9$ we see we are restricted to the narrow range $m_{\chi}\sim(10^{7}-10^{8})\,GeV$ for Hubble scales at the end of inflation $H_{\rm end}\sim(10^{8}-10^{9})\,GeV$. On the other hand if the conductivity redshifts like Hubble ($\alpha=2$) or faster then we are restricted to much larger masses in the range $m_{\chi}\sim(10^{8}-10^{10})\,GeV$ regardless of the value of $\xi_{\rm end}$ and $\epsilon_{R}$. In this case, the dark matter is generated almost entirely at the end of inflation.
In summary we can have dark charged particles as the dark matter for masses as light as $\sim 10\,MeV$ if the conductivity is constant after inflation and as heavy as $m_{\chi}\sim 10^{10}\,GeV$. This mechanism generically prefers larger Hubble scales at the end of inflation in the range $H_{\rm end}\sim(10^{8}-10^{14})\,GeV$. In the end we are left with dark electrons (scalar or fermion) as the dark matter under the influence of a feeble dark QED force mediated by a massless (or very light) dark photon.
We can also examine the parameter space independently of the mechanism for generating the dark electric field by plotting contours of $\Omega_{\chi}/\Omega_{\rm CDM}=1$ as a function of the ratios $g_{D}E/H_{\rm end}^{2}$ vs. $m_{\chi}/H_{\rm end}$. We show these contours in Fig. 2 for $\epsilon_{R}=0.75$ and only the $\alpha\gg 2$ case where all of the production occurs during inflation and thus is independent of the unknown conductivity after reheating. This also allows for comparison to the parameter space in the strong force Bastero-Gil et al. (2023) regime ($g_{D}E>H^{2}$) where the Schwinger production necessarily occurs only during inflation and requires $\epsilon_{R}\gtrsim 0.5$. We see that while in the strong force regime we are limited to $m_{\chi}\gtrsim H_{\rm end}\sim 10^{14}\,GeV$, the weak force case can accommodate Hubble scales down to $H_{\rm end}\sim 10^{8}\,GeV$ and dark matter masses in the range $m_{\chi}\simeq(10^{8}-10^{10})\,GeV$. These results hold for any mechanism which can generate a dark electric field during inflation which leads to Schwinger production of dark electron pairs.
Figure 2: Contours of $\Omega_{\chi}/\Omega_{\rm CDM}=1$ for $\epsilon_{R}=0.75$ and $\alpha\gg 2$ as a function of $g_{D}E/H_{\rm end}^{2}$ vs. $m_{\chi}/H_{\rm end}$. Above the contour lines we have overabundance, below them underabundance. The light blue region is excluded by the non-relativistic constraint in Eq. (49).
The dark QED sector contained in this mechanism could have a number of interesting phenomenological possibilities. The dark photon could potentially mix with the visible photon leading to a mili-charged dark matter sector with interesting phenomenology Berlin et al. (2023). It could also have interesting implications for structure formation as the dark coulomb interaction must be screened at length scales of roughly the size of cluster of galaxies or it would be stronger than the gravitational interaction. Some screening occurs due to the fact that the dark photon will inevitably get an effective mass (plasma frequency) through the interaction with the dark electron plasma. We can estimate the effective mass at redshift around 100 as $\omega_{p}(T_{*})=g_{D}\sqrt{\rho_{\chi}(T_{*})}/m_{\chi}$ taking $T_{*}\simeq 100\ T_{0}$. This gives $\omega_{p}(T_{*})\sim g_{D}10^{4}\ {\rm Mpc}^{-1}$ for $m_{\chi}\sim 10^{14}$ GeV which implies that for a dark gauge coupling $\gtrsim 10^{-4}$ there should be enough screening. We leave a detailed study of the impact of our scenario on structure formation to future work. One should also worry about constraints from the observed triaxality of dark matter halos Agrawal et al. (2017) and from plasma instabilities Lasenby (2020), but because the dark gauge coupling in the mechanism considered here is so small we should be safe from these bounds. Finally, assisted Schwinger production Siemonsen et al. (2023) due to accelerated dark electrons could potentially destroy the dark electric field, but since the dark electrons reach terminal velocity (see Eq. (47)) so quickly we expect this to also have little effect. We leave detailed investigations of these questions to future work.

### v2

@@ III.1 Constraints on dark gauge coupling
As discussed above, in principle there are severe constraints on $g_{D}$ coming from structure formation and plasma insatiabilities Lasenby (2020), but these can be avoided if the dark vector has a mass $m_{A}\gtrsim 10^{-10}\,eV$. Assisted Schwinger production Siemonsen et al. (2023) due to accelerated dark electrons could potentially destroy the dark electric field and constrain $g_{D}$, but since the dark electrons reach terminal velocity (see Eq. (34)) so quickly during inflation and are not accelerated afterwards (see Appendix) we expect this to also have little effect. When examining the dark matter parameter space below we thus only impose on $g_{D}$ the constraints in Eq. (37)-Eq. (40).
III.2 Dark matter parameter space
@@
With the various energy densities today in hand we can examine the parameter space in which the observed dark matter relic abundance is reproduced. Taking the observed energy density of cold dark matter today Ade et al. (2016) to be $\rho_{\rm CDM}=9.6\times 10^{-48}\ {\rm GeV}^{4}$, the relic abundances of the dark vector and dark electrons are given by,
$\frac{\Omega_{A}}{\Omega_{\text{CDM}}}$ $=$ $\frac{\rho_{A}(T_{0})}{\rho_{\text{CDM}}},\quad\frac{\Omega_{\chi}}{\Omega_{\text{CDM}}}=\frac{\rho_{\chi}(T_{0})}{\rho_{\text{CDM}}}\,,$ (41)
where $\rho_{A}(T_{0})$ and $\rho_{\chi}(T_{0})$ are given in Eq. (33) and Eq. (36) respectively. With these estimates for the relic abundance, we can go on to obtain the regions of parameter space for a viable dark matter sector. In addition to the constraints on $g_{D}$ discussed above, there are additional constraints on the model parameters which must be satisfied for consistency of the mechanism.
The first is that we must also require the hierarchy in energy densities at reheating in Eq. (27). Imposing that the visible radiation has energy density lower than the inflaton at the end of inflation ($\rho_{R}\ll\rho^{\rm{end}}_{I}$) imposes $\epsilon_{R}<1$. To ensure the Universe is radiation dominated until matter radiation equality requires that $\rho_{R}>\rho^{\rm{end}}_{E}$ and translates into an upper bound on $H_{\rm end}$,
$H_{\rm end}<\frac{\sqrt{6}\,\epsilon_{R}^{2}M_{\rm Pl}}{(E/H_{\rm end}^{2})}\,.$ (42)
So we see that for very large electric fields or reheating temperatures which are low $(\epsilon_{R}\ll 1)$, the upper bound on the Hubble scale at the end of inflation becomes very stringent. We impose the stronger between this upper bound and the absolute astrophysical upper bound of $H_{\rm end}\lesssim 10^{13}$ GeV which must be smaller than the inflationary scale early during inflation. We must also ensure no backreaction on the inflaton dynamics from production of the dark electric field. This requires $\rho^{\rm{end}}_{E}<\rho^{\rm{end}}_{I}$ which restricts the size of the electric field to Bastero-Gil et al. (2019, 2022),
$1<\frac{E}{H_{\rm end}^{2}}<10^{9}.$ (43)
Note that for the vector production mechanism discussed above, this corresponds to $1.6\lesssim\xi_{\rm end}\lesssim 9$. The dark electron mass is constrained from below by requiring that $\chi$ becomes non-relativistic before matter radiation equality, thus behaving like cold dark matter. In practice the no-thermalization condition imposes stronger lower bounds on $m_{\chi}$. There are also robust constraints coming from astrophysical probes which in the case of fermions can lead to a stronger and absolute lower bound of $\mathcal{O}(100\ {\rm eV})$ Di Paolo et al. (2018); Alvey et al. (2021) for fermion dark matter.
We focus on the light mass case where $m_{\chi}<H_{\rm end}$ while the large mass case $m_{\chi}>H_{\rm end}$ was examined in Bastero-Gil et al. (2023) to which we refer the interested reader. We assume the dark photon mass is larger than $m_{A}\gtrsim 10^{-10}$ eV to avoid severe constraints on $g_{D}$ coming from plasma instabilities Lasenby (2020), but small enough that it contributes negligibly to the dark matter relic abundance.
III.2.1 Weak dark electric force $\lambda<1$
We first consider the weak electric force regime where $\lambda<1$. Note this case is not possible in Minkowski space which requires a strong electric force and is a unique feature in de-sitter space Banyeres et al. (2018). In this regime the upper bound on $g_{D}$ coming from imposing $\lambda<1$ is a stronger constraint than both the no thermalization and no Schwinger backreaction conditions. In Fig. 1 we show contours of $\Omega_{\chi}/\Omega_{\rm{CDM}}=1$ given in Eq. (41) as a function of $m_{\chi}$ vs. $H_{\rm end}$ for various values of $E/H_{\rm end}^{2}$ and $\epsilon_{R}$ as indicated. For the conductivity in the $\lambda<<1$ regime we take the conformal limit $m_{\chi}=0$ (and $\xi_{R}=1/6$ in the scalar case defined in Eq. (2)) leading to Eq. (19). Note since we are in the conformal limit, purely gravitational production is negligible Ford (1987); Kuzmin and Tkachev (1999).
Figure 1: Contours of $\Omega_{\chi}/\Omega_{\rm{CDM}}=1$ given in Eq. (41) as a function of $m_{\chi}$ vs. $H_{\rm end}$ in the weak electric force regime where $\lambda<1$ for various values of $E/H_{\rm end}^{2}$, $\epsilon_{R}$, and $g_{D}$ as indicated. See text for more information.
In the top left plot we show contours for $E/H_{\rm end}^{2}=1$ and $\epsilon_{R}=0.5$. For the dark electric field production mechanism discussed above, $E/H_{\rm end}^{2}=1$ corresponds to $\xi_{\rm end}\simeq 1.6$ (see Eq. (21)). Note that for $\epsilon_{R}=0.5$ this range of Hubble scales corresponds to reheating temperatures between $T_{\rm RH}\sim(10^{9}-10^{15})$ GeV (see Eq. (23)). For $E/H_{\rm end}^{2}=1$ the maximum value of $g_{D}$ allowed by imposing $\lambda<1$ is given by $g_{D}^{\rm{max}}=1$ (black) for the which the dark matter relic abundance can be reproduced over a very large range of masses from $m_{\chi}\sim 10^{-1}$ GeV to $m_{\chi}\sim 10^{8}$ GeV. For this parameter point, dark gauge couplings as small as $g_{D}\simeq 10^{-6}$ can reproduce the dark matter relic abundance for dark matter masses $m_{\chi}\sim 10^{12}$ GeV and $H_{\rm end}\sim 10^{13}$ GeV. Much smaller values of $g_{D}$ would require $H_{\rm end}\gtrsim 10^{13}$ GeV which is excluded by astrophysical measurements.
On the top right we show contours for $E/H_{\rm end}^{2}=10^{9}$ ($\xi_{\rm end}\simeq 9$) which is the maximum allowed size for the electric field before backreaction affects the inflaton dynamics Barnaby et al. (2012); Meerburg and Pajer (2013). For $E/H_{\rm end}^{2}=10^{9}$ the maximum value of $g_{D}$ allowed by imposing $\lambda<1$ is $g_{D}^{\rm{max}}=10^{-9}$ (black). We see the much larger dark electric fields push us to lower Hubble scales in order to ensure $\rho^{\rm{end}}_{E}<\rho_{R}$. Furthermore, since the $\lambda<1$ condition forces us to very small dark gauge couplings, it is more difficult to reproduce the relic abundance. Even so, we see there is still a small range of masses $m_{\chi}\sim(10^{6}-10^{8})$ GeV and Hubble scales $m_{\chi}\sim(10^{8}-10^{9})$ GeV for which the relic abundance can be reproduced with $g_{D}\sim(10^{-9}-10^{-10})$.
On the bottom left we show contours for $E/H_{\rm end}^{2}=10^{4}$ ($\xi_{\rm end}\simeq 5$) with $\epsilon=0.5$. Now the maximum value of $g_{D}$ allowed by imposing $\lambda<1$ is $g_{D}^{\rm{max}}=10^{-4}$ (black). Overall the parameter space is similar to the top left figure with $E/H_{\rm end}^{2}=1$ and $\epsilon=0.5$ except now we are pushed to smaller gauge couplings which can be as small as $g_{D}\simeq 10^{-8}$ for $m_{\chi}\sim 10^{9}$ GeV.
On the bottom right we show contours for $E/H_{\rm end}^{2}=10^{4}$ with $\epsilon_{R}=10^{-5}$. In this case we have lower reheating temperatures in the range $T_{\rm RH}\sim(10^{4}-10^{10})$ GeV. This forces us to smaller Hubble scales in order to ensure $\rho^{\rm{end}}_{E}<\rho_{R}$. This in turn lowers $m_{\chi}$ in order to ensure $m_{\chi}<H_{\rm end}$. We see that for dark gauge couplings in the range $g_{D}\simeq(10^{-4}-10^{-7})$ the relic abundance can be reproduced for dark matter masses in the range $m_{\chi}\sim(10^{-2}-10^{5})$ GeV and Hubble scales at the end of inflation in the range $H_{\rm end}\simeq(100-10^{5})$ GeV.
To summarize, in the weak force regime with $\lambda<1$, inflationary dark Schwinger production can reproduce the dark matter relic abundance for dark electron masses in the range $m_{\chi}\simeq(10^{-1}-10^{12})$ GeV and gauge couplings between $g_{D}\simeq 10^{-10}-1$ for Hubble scales at the end of inflation in the range $H_{\rm end}\simeq(100-10^{13})$ GeV.
III.2.2 Strong dark electric force $\lambda>1$
In Fig. 2 we show contours of $\Omega_{\chi}/\Omega_{\rm{CDM}}=1$ in the strong electric force regime where $\lambda>1$. In this regime the Schwinger conductivity is the same as for flat space Banyeres et al. (2018) and is the same for both fermions and scalars (up to a factor of 2). The $\lambda>1$ condition now imposes a lower bound on $g_{D}$ while an upper bound (blue shaded) comes from requiring that the dark sector does not thermalize with itself (see Eq. (37)). There is also an upper bound from requiring no Schwinger backreaction (red shaded), but it is weaker than requiring no thermalization. In principle the parameter space between these two bounds could also be viable dark matter parameter space, but one would need to track the thermal history after inflation. This involves other strong dark QED processes Arvanitaki et al. (2021) in order to compute the proportion of dark electrons to photons in the final relic abundance and is beyond the scope of this work. We show this region for illustrative purposes, but when considering the viable dark matter parameter space presented here we take the no-thermalization condition to give the upper bound on $g_{D}$.
Figure 2: Contours of $\Omega_{\chi}/\Omega_{\rm{CDM}}=1$ given in Eq. (41) as a function of $m_{\chi}$ vs. $H_{\rm end}$ in the strong electric force regime where $\lambda>1$ for various values of $E/H_{\rm end}^{2}$, $\epsilon_{R}$, and $g_{D}$ as indicated. See text for more information.
In the top left plot we show contours again for $E/H_{\rm end}^{2}=1$ and $\epsilon_{R}=0.5$. We see in this case there is a narrow band between $g_{D}\simeq 1-2$ for which the relic abundance can be reproduced. Along this band however the relic abundance can be reproduced over a large range of dark matter masses from $m_{\chi}\simeq(1-10^{8})$ GeV for $10^{8}\lesssim H_{\rm end}/{\rm GeV}\lesssim 10^{13}$.
On the top right we show contours for $E/H_{\rm end}^{2}=10^{9}$ and $\epsilon_{R}=0.5$. This case gives a large parameter space in which the relic abundance can be reproduced for dark gauge coupling in the range $g_{D}\simeq 10^{-9}-0.002$ and $100\lesssim H_{\rm end}\lesssim 10^{9}$ GeV. For $g_{D}$ near the thermalization bound of $g_{D}\approx 0.0002$ a dark charged scalar as light as $m_{\chi}\sim 0.1$ eV can give the correct relic abundance. In the fermion case astrophysical constraints impose an absolute lower bound of $100$ eV (dashed grey line).
On the bottom left we show contours for $E/H_{\rm end}^{2}=10^{4}$ with $\epsilon=0.5$. We see dark electrons as light as $m_{\chi}\sim 100$ eV can give the correct relic abundance, but now for dark gauge couplings near the thermalization bound of $g_{D}\simeq 0.04$ and Hubble scales $H_{\rm end}\sim 10^{13}$ GeV.
Finally, on the bottom right we show contours for $E/H_{\rm end}^{2}=10^{4}$ with $\epsilon_{R}=10^{-5}$. Again the lower reheating temperatures force us to smaller Hubble scales in order to ensure $\rho^{\rm{end}}_{E}<\rho_{R}$. In the end the relic abundance can still be reproduced for a narrow band $g_{D}\simeq 10^{-4}-10^{-3}$ and $H_{\rm end}\simeq(100-10^{5})$ GeV leading to a dark matter mass in the range $m_{\chi}\simeq(10^{-4}-100)$ GeV.
In summary, in the $\lambda>1$ strong force regime we can have dark charged scalars as the dark matter for masses as light as $\sim 0.1$ eV and $100$ eV for fermions and as heavy as $m_{\chi}\sim 10^{12}$ GeV in both cases. The dark gauge coupling can span the range $110^{-9}-$ 1 while the Hubble scale at the end of inflation can span essentially the entire range allowed by astrophysical bounds.